I AM A PARENT TRYING TO HELP MY DAUGHTER DO THIS PROBLEMS I NEVER HAD THIS MAYBE YOU COULD SHOW ME HOW SO I CAN HELP HER.


SOLVE LINER COMBINATION TO SOLVE THE SYSTEM


2X -3Y = -16
X + 3Y = 10



3X + Y = -8
-3 + 4Y = -2

2X-Y = 1
2X + 5Y = -5


THIS SAY USE ANY METHOD TO SOLVE

-2X + 3Y = 14
X-4Y = -12


IT BEEN A LONG TIME I DON'T EVEN REMEMBER HAVING THIS

Found 2 solutions by venugopalramana, fractalier:
Answer by venugopalramana(3286)

you gave me a site to go on and i can't get on it i am trying to see how to do these because i am out of school right now and hve no one to show me
use linear combinations to solve the system linera equation
THESE ARE ALL SIMILAR PROBLEMS.I SHALL SHOW ONE AND YOU MAY TRY THE OTHERS.
1. x + 2 =5....................EQN.I
3x - 2y =7.....................EQN.II
1.THERE ARE 2 UNKNOWNS AND 2 EQNS.
2.IF THERE IS 1 UNKNOWN AND 1 EQN.WE CAN SOLVE IT
3.SO WE TRY TO REMOVE ONE UNKNOWN BY ELIMINATION.
EQN.I GIVES X=5-2=3
EQN.II ..3*3-2Y=7
2Y=9-7=2
Y=1
2. x + y = 1......................III
2x -3y = 12........................IV
EQN.I*2-EQN.II
2X+2Y-2X+3Y=2-12=-10
5Y=-10
Y=-2
PUT THIS IN EQN.III
X-2=1
X=3....ETC..
3. x -y = -4
x + 2y = 5

Answer by fractalier(6550)

Okay, let me show you...the idea of linear combination is to combine equations in such a way (by adding or subtracting them) that eliminates one of the variables...by example...we can add these equations
2X -3Y = -16
X + 3Y = 10
and get
3X = -6
and X = -2
Notice how the 3Y and -3Y cancelled each other out when we added them...
From this X = -2, we plug in to either equation to find Y = 4
Second one...same thing...
3X + Y = -8
-3X + 4Y = -2
we add them to get
5Y = -10
and Y = -2...plug in and get X = -2
On the third one we subtract equations because if we added them, nothing would cancel...
2X - Y = 1
-(2X + 5Y = -5)
and get
-6Y = 6 and
Y = -1 and thus X = 0
The last one is a bit more complicated s you have to multiply the bottom equation by two before adding them, so from
-2X + 3Y = 14
X - 4Y = -12
we have
-2X + 3Y = 14
2X - 8Y = -24
and we add giving us
-5Y = -10
Y = 2 and x = -4
I have five kids of my own...I know how you feel...
Buy me a beer next time you see me...